HELP
Calculate the work done by the force field F on an object moving along a curve from P(-5, 1) to Q(7,2). F(x, y) =
Answer #1

Fexry) - de P (-511) to 0 (712) F(y) = 27 2 - 2 4 5 =pita] =-22 P= 27 2 OP Do y = - 3-92 ye y? E is conservative oy Spax af a